# Do Large Language Models Think Like the Brain? Sentence-Level Evidence from fMRI and Hierarchical Embeddings

### v1

@@ 4.3 Left-right Hemispheric Asymmetry
To examine the relationship between hemispheric asymmetry and model performance, Figure 8(c) shows correlations between LH-RH differences and performance metrics. Among the six ROIs analyzed, IFG and MFG lateralization were potentially linked to performance, showing positive relationships (Pearson correlation: IFG: $r=0.54$, $p=0.055$; MFG: $r=0.50$, $p=0.084$). No significant correlations were found for other regions(Pearson correlation: AntTemp: $r=0.02$, $p=0.941$; PostTemp: $r=0.09$, $p=0.770$; AngG: $r=0.09$, $p=0.687$; IFGorb: $r=-0.06$, $p=0.845$). The regulatory role of the prefrontal cortex over distributed brain regions may underlie this phenomenon Badre and Nee (2018), where its lateralized functional specialization could enhance top-down coordination of cognitive resources, mirroring the efficiency optimization observed in computational models.
5 Discussion
Understanding how LLMs align with human brain activity during sentence processing provides critical insights into both artificial and biological language systems Mahowald et al. (2024); Zhou et al. (2024).
From the perspective of brain decoding in sentence tasks, our model also exhibits higher activation correlations in the middle layers compared to the final layer, which is consistent with the findings of Mischler et al. (2024). Importantly, our exploration using fMRI data complements their approach based on intracranial electrode recordings and MEG Zhou et al. (2024). At the level of sentence processing, Yu et al. (2024) investigated the process from the perspective of BERT, whereas we not only consider BERT but also compare it with larger LLMs (6.7B–9B parameters). We further find that instruction-tuned models outperform base models in both correlation and sentence comprehension ability, with marginally significant improvements in correlation. This observation is in line with Ren et al. (2025), which also reported that instruction tuning can significantly enhance LLM-brain similarity. In addition, while previous studies such as Bonnasse-Gahot and Pallier (2024) primarily focus on the relationship between model size and brain correlation, our work shifts the emphasis to the model’s comprehension ability. Specifically, within the parameter range of 6.7B to 9B, we reveal not only a significant positive association between comprehension ability and neural similarity, but also identify a scaling law: as the comprehension ability of language models increases, their neural alignment with brain activity systematically and predictably improves.
Left-right hemispheric asymmetry is a fundamental organizational principle of the human brain, shaping the specialization of language and cognitive functions. The left-hemispheric dominance in IFG and PostTemp aligns with their established roles in core language functions Hu et al. (2023), supporting syntactic encoding and semantic integration. In contrast, right-hemispheric engagement in MFG and AntTemp may reflect specialized processes for metaphor comprehension, contextual integration Japee et al. (2015), and multimodal semantic representations. The positive trends between leftward IFG/MFG asymmetry and model performance suggest that enhanced neural specialization in these regions may reduce interhemispheric interference and improve processing efficiency Cai et al. (2013); Friedman and Robbins (2022). Stronger left IFG activation has been previously linked to superior language capabilities Gotts et al. (2013); Arredondo et al. (2019). Non-significant correlations in other regions may indicate bilateral coordination or domain-general processes, such as the AngG’s role in the default mode network and non-lateralized semantic integration Kuhnke et al. (2023).
These findings highlight that LLM processing engages both conserved left-hemispheric language systems and right-hemispheric networks supporting higher-order cognitive demands. The functional specificity of hemispheric asymmetries underscores the need to disentangle distinct neural contributions to LLM performance, rather than attributing it to a singular mechanism.
6 Conclusion
This study bridges LLMs and human brain activity by introducing a custom sentence understanding task paired with naturalistic fMRI data. Our evaluation of 14 open LLMs reveals that instruction tuning substantially enhances both task performance and alignment with neural responses, demonstrating statistically significant correlations. Additionally, we observe that the performance of the intermediate layers surpasses that of the final layer within the LLMs. Moreover, we identify clear patterns of left-hemispheric dominance in key language regions, suggesting efficient, specialized processing, alongside function-specific asymmetries that reflect differentiated neural roles in supporting language comprehension. Together, these findings demonstrate that model architecture and training objectives shape the degree of brain alignment, and highlight the value of integrating cognitively grounded tasks with neuroimaging to better understand the neural basis of language and guide future model development.
Limitations
This study was limited to neurotypical participants, without inclusion of individuals with language impairments or neurodevelopmental diversity. As a result, the clinical relevance of the findings is constrained, as atypical language processing patterns in populations such as individuals with aphasia or dyslexia were not captured. Additionally, the participant sample was culturally homogeneous, consisting primarily of native Chinese speakers, which limits the generalizability of the results to cross-linguistic or bilingual populations. These constraints may affect the applicability of our conclusions, particularly in clinical or cross-cultural contexts, and highlight the need for more inclusive sampling in future research.
Ethics Statements
The fMRI data we used in the current study are from a publicly available resource Li et al. (2022). The current study was reviewed by the IRB of Jiangsu Normal University. The study itself does not raise any ethical concerns.
Acknowledgments
The authors acknowledge the use of ChatGPT exclusively for grammatical checks in the final manuscript.
Appendix A Details on Regions of Interest (ROIs)
The language network occupies a central role in the human brain, comprising the lateral frontal cortex (encompassing the inferior frontal gyrus and middle frontal gyrus) and the lateral temporal cortex (involving the superior and middle temporal gyri extending from the temporal pole to the posterior temporal lobe). In the majority of individuals, these regions demonstrate functional lateralization, with the left hemisphere exhibiting stronger and more spatially extensive activation during linguistic processing. Notably, emerging evidence highlights the right hemisphere’s crucial contributions to language information processing, particularly manifesting distinct patterns of neural activation during higher-order linguistic functions such as pragmatic inference, metaphor comprehension, and affective prosody processing.
The network demonstrates remarkable functional specificity: The sentence > nonwords sequence contrast paradigm effectively isolates selective responses to linguistic information (particularly syntactic structuring and semantic integration) while suppressing neural engagement in non-linguistic cognitive tasks such as arithmetic reasoning, working memory, and spatial processing. This functional decoupling from the multiple-demand network Fedorenko et al. (2024) ensures focused analysis of core discourse comprehension processes, eliminating confounding effects from perceptual encoding and executive functions. Such specificity proves particularly advantageous for analyzing texts like The Little Prince that demand deep syntactic-semantic integration due to their complex metaphorical structure and narrative architecture.
Second, the network’s neural representations exhibit cross-modal and cross-linguistic universality: Large-scale neuroimaging data encompassing 45 languages Malik-Moraleda et al. (2022) reveal stable topological organization across visual/auditory modalities and comprehension/production tasks. This domain-general consistency establishes it as an ideal framework for exploring fundamental mechanisms of human language capacity, while providing theoretical grounding for analyzing multilingual translation processing mechanisms.
Methodologically, Fedorenko’s team addressed inter-individual functional-anatomical variability through probabilistic functional localization derived from fMRI data of 800 participants Lipkin et al. (2022). Compared to traditional anatomical ROI definitions, this approach significantly enhances functional region reproducibility (inter-subject correlation coefficients >0.85), ensuring superior ecological validity. Empirical evidence confirms the network’s heightened sensitivity to incremental parsing and narrative integration during natural language processing Fedorenko et al. (2016), effectively capturing neural signatures of advanced semantic operations like metaphor comprehension. Notably, geometric isomorphism exists between the network’s neural representations and deep-layer embeddings in Transformer architectures Tuckute et al. (2024). This representational correspondence provides crucial theoretical support for comparative analysis of processing mechanisms between biological and artificial neural networks within a unified functional framework.
Region Name Abbreviation Functional Domain
Left Hemisphere Inferior frontal gyrus LH_IFG
• Higher-order language processing
• Lexical retrieval, syntax integration
• Distributed syntactic processing
Left Hemisphere Inferior frontal gyrus orbital LH_IFGorb
• Higher-order language processing
• Lexical/sentence meaning and structure
• Reading
Left Hemisphere Middle frontal gyrus LH_MFG
• Language processing
• Linguistic working memory
Left Hemisphere Anterior temporal LH_AntTemp
• Semantic composition
• Construction of complex meaning
• Longer temporal integration windows
Left Hemisphere Posterior temporal LH_PostTemp
• Core language computation “hub”
• Lexical retrieval and basic semantic integration
• Shorter temporal processing windows
Left Hemisphere Angular Gyrus LH_AngG
• Semantic integration
• Discourse comprehension
• Interface with social cognition systems
Right Hemisphere Inferior frontal gyrus RH_IFG
• Semantic control in sentence/narrative processing
• Supports language comprehension under cognitive load
Right Hemisphere Inferior frontal gyrus orbital RH_IFGorb
• Supports language processing under difficulty
• May engage in demanding tasks
Right Hemisphere Middle frontal gyrus RH_MFG
• Supports language processing under difficulty
• May engage in demanding tasks
Right Hemisphere Anterior temporal RH_AntTemp
• Processes social content in language
• Supports language under difficulty
• Overlaps with ToM-related areas
Right Hemisphere Posterior temporal RH_PostTemp
• Processes social content in language
• Integrates contextual information
• Supports language under difficulty
Right Hemisphere Angular Gyrus RH_AngG
• Supports language processing under difficulty
• Sensitive to non-language ToM tasks
• May be recruited for socially driven reasoning
Table 2: Brain regions, their abbreviations, and associated functional domains in language processing.
Abbreviations: LH – Left Hemisphere; RH – Right Hemisphere; IFG – Inferior Frontal Gyrus; MFG – Middle Frontal Gyrus; AntTemp – Anterior Temporal; PostTemp – Posterior Temporal; AngG – Angular Gyrus.
Appendix B Details of General Linear Model (GLM)
The neuroimaging analysis employs a mass-univariate General Linear Model (GLM) framework to estimate task-evoked neural activity from blood-oxygen-level-dependent (BOLD) signals. The core model is expressed as:
where $\mathbf{Y}\in\mathbbm{R}^{T\times V}$ denotes the preprocessed BOLD time series across $T$ time points (TRs) and $V$ voxels, $\mathbf{X}\in\mathbbm{R}^{T\times P}$ the design matrix encoding experimental conditions and nuisance regressors, $\boldsymbol{\beta}\in\mathbbm{R}^{P\times V}$ the unknown regression coefficients, and $\boldsymbol{\epsilon}$ the residual error with homoscedastic variance $\sigma^{2}$.
The design matrix $\mathbf{X}$ is constructed through the convolution of trial onsets with a canonical hemodynamic response function (HRF). For each experimental condition $j$ with $K_{j}$ events occurring at times $\{t_{\text{onset}}^{(j,k)}\}_{k=1}^{K_{j}}$, the corresponding regressor $\mathbf{x}_{j}$ is generated as:
where $\text{HRF}(\cdot)$ represents the double-gamma function from SPM12, and $\{\mathbf{n}_{m}(t)\}_{m=1}^{M}$ capture nuisance confounds (e.g., motion parameters, drift terms) weighted by coefficients $\gamma_{m}$.
In the present study, neural responses elicited by complex linguistic stimuli were estimated using a single-trial modeling framework based on the Least-Squares Separate (LS-S) method Mumford et al. (2014). This analytical strategy enables sentence-level independent regression, thus facilitating a more precise dissociation of blood-oxygen-level-dependent (BOLD) signals attributable to individual trials. By modeling each target sentence separately while controlling for variance introduced by non-target trials, this approach substantially mitigates the issue of condition collinearity—a pervasive limitation in conventional block-based or condition-averaged designs. In contrast to the alternative Least-Squares All (LS-A) technique, which models all trial events concurrently, the LS-S method retains inter-trial temporal variability and improves sensitivity to transient, event-related cognitive processes through distinct parameter estimation for each trial. This increased granularity in modeling improves the reliability and interpretability of neural activation patterns associated with sentence-level processing under naturalistic language input conditions.
Parameter estimation via ordinary Least Squares-Separate(LS-S) yields voxel-wise coefficient maps:
followed by computation of $t$-statistics for hypothesis testing on linear contrasts $\mathbf{c}\in\mathbbm{R}^{P}$:
where $t_{v}$ quantifies the standardized effect size for voxel $v$, and $\hat{\sigma}_{v}^{2}$ estimates the residual variance. This formulation supports inference on condition-specific activation patterns while controlling for physiological noise and scanner drift.
Appendix C Details of Cross-Validated Ridge Regression for Layer-wise Encoding Analysis
The neurocognitive encoding pipeline utilizes regularized regression to quantify the semantic representations associated with specific layers of a deep neural network in brain activity patterns derived from functional magnetic resonance imaging (fMRI). This framework bridges computational neuroscience and artificial intelligence, enabling a systematic analysis of the relationship between brain responses and representations derived from deep learning models. Specifically, given the fMRI response vectors $\mathbf{y}\in\mathbbm{R}^{T}$ from a region of interest (ROI) $r$ and deep learning embeddings $\mathbf{X}\in\mathbbm{R}^{T\times L\times D}$ (where $T$ is the number of time points, $L$ is the number of neural network layers, and $D$ is the embedding dimensionality), the model’s layer-wise predictive performance is computed as follows:
$\mathcal{L}(\mathbf{X}^{(l)},\mathbf{y})=\frac{1}{K}\sum_{k=1}^{K}\rho\left(\mathbf{y}_{\text{test}}^{(k)},\hat{\mathbf{y}}_{\text{test}}^{(k)}\right)$ (13)
Here, $\rho(\cdot)$ refers to the Pearson correlation coefficient, a common metric to assess the similarity between actual fMRI responses $\mathbf{y}_{\text{test}}^{(k)}$ and predicted responses $\hat{\mathbf{y}}_{\text{test}}^{(k)}$ from the ridge regression model. Ridge regression is applied to compute $\hat{\mathbf{y}}$, regularizing the regression coefficients $\boldsymbol{\beta}$ to prevent overfitting in high-dimensional datasets:
$\hat{\boldsymbol{\beta}}^{(l)}=\arg\min_{\boldsymbol{\beta}}\left\|\mathbf{X}_{\text{train}}^{(l)}\boldsymbol{\beta}-\mathbf{y}_{\text{train}}\right\|^{2}+\alpha\|\boldsymbol{\beta}\|^{2}$ (14)
In this equation, $\mathbf{X}_{\text{train}}^{(l)}$ and $\mathbf{y}_{\text{train}}$ represent the training inputs and fMRI targets, respectively, while the regularization parameter $\alpha$ controls the trade-off between data fidelity and model complexity by penalizing large values of $\boldsymbol{\beta}$. This ensures the model is robust to noise and can generalize better to unseen testing data.
Feature Normalization.
To ensure numerical stability and compatibility between fMRI responses $\mathbf{y}$ and model embeddings $\mathbf{X}$, all variables are z-scored prior to regression:
$\mathbf{y}\rightarrow\frac{\mathbf{y}-\mu_{y}}{\sigma_{y}},\quad\mathbf{X}^{(l)}\rightarrow\frac{\mathbf{X}^{(l)}-\boldsymbol{\mu}_{X}}{\boldsymbol{\sigma}_{X}}$ (15)
Here, $\mu_{y}$ and $\sigma_{y}$ are the mean and standard deviation of the fMRI signals across time, ensuring that $\mathbf{y}$ has zero mean and unit variance. Similarly, $\boldsymbol{\mu}_{X}$ and $\boldsymbol{\sigma}_{X}$ represent the channel-wise mean and standard deviation of the neural embeddings $\mathbf{X}^{(l)}$. This normalization step ensures that all features (both neural and brain signals) are on comparable scales, improving the stability and interpretability of the regression weights.
Layer-wise Regression Analysis.
For each layer $l$ of the neural network ($l\in\{1,...,L\}$), we evaluate the performance of the model in predicting the fMRI responses. Ridge regression is applied independently on each layer’s embeddings $\mathbf{X}^{(l)}$ to fit the training data. The prediction quality for that layer is quantified as the average Pearson correlation across $K$ cross-validation folds:
Here: - $\hat{\boldsymbol{\beta}}^{(l,k)}$ is computed from the training subset $(\mathbf{X}_{\text{train}}^{(l,k)},\mathbf{y}_{\text{train}}^{(k)})$ using the closed-form solution:
This equation balances data fitting ($\mathbf{X}^{\top}\mathbf{X}$) and regularization (the identity matrix $\mathbf{I}$ scaled by $\alpha$). - The hyperparameter $\alpha$ is selected through grid search (e.g., $\alpha\in\{10^{-3},10^{-2},...,10^{1}\}$) optimized within a nested cross-validation loop to maximize the prediction accuracy.
This process produces a layer-wise correlation metric $\rho_{l}$, which reflects how well the embeddings at that neural layer predict ROI-specific brain responses.
Hierarchical Computational Parallelization.
Given that layer-wise analysis needs to be performed across multiple subjects, ROIs, and neural network layers, the computational pipeline is structured as a nested series of parallel loops:
$\underbrace{\text{Subjects}}_{\text{Outer loop}}\times\underbrace{\text{ROIs}}_{\text{Middle loop}}\times\underbrace{\text{Layers}}_{\text{Inner loop}}\rightarrow\mathcal{O}(SRL)\text{ models}$ (18)
Here, $S$ represents the number of subjects, $R$ represents the number of ROIs analyzed, and $L$ is the total number of neural network layers. In practice, this results in training and evaluating an extensive number of ridge regression models ($\mathcal{O}(SRL)$), thereby motivating an efficient parallel implementation.
Appendix D Input Perturbations
We show in the following the detailed input perturbations we employed for the LLMs sentence performance tasks we conducted in §4.1.
  • Option A is the correct English translation of the Chinese sentence.
  • Option B is constructed by scrambling the word order of Option A, thereby disrupting the syntactic structure while preserving the lexical content.
  • Option C is formed by substituting certain words in Option A with their synonyms or words of the same part of speech, altering the surface semantics.
  • Option D is generated by transforming the sentence structure of Option A, such as changing from active to passive voice, while maintaining the original meaning.
  • Option E is created by inserting or deleting information in Option A, thus introducing semantic noise or omitting key details.
Appendix E Details on the LLMs
We present the detailed information of the 14 LLMs in Table 3. During the experiments, we used the transformers Wolf et al. (2020) and pytorch Paszke et al. (2019) library for training the models. All experiments were conducted on an NVIDIA® A100 80 GB RAM GPU.
Scaling Model Layers
~110M Bert-base-uncased 12
~7B Qwen2.5-7B 28
Qwen2.5-7B-Instruct 28
Mistral-7B-v0.3 32
Mistral-7B-Instruct-v0.3 32
Baichuan2-7B-Chat 32
DeepSeek-R1-Distill-Qwen-7B 32
opt-6.7b 32
~8B Llama-3.1-8B 32
Llama-3.1-8B-Instruct 32
~9B gemma-2-9b 28
gemma-2-9b-it 28
Glm-4-9b-chat-hf 32
Glm-4-9b-hf 32
Table 3: Layer numbers of transformer-based models across different parameter scaling levels. Model names link to their Hugging Face repositories.
Appendix F Additional Experimental Results
Model Performance and Activation Correlation.
To comprehensively assess the correspondence between computational models and brain activity, we compared model performance across multiple brain regions and layers, as addition to §4.2. Figure 9 displays the distribution of correlation coefficients for 12 bilateral regions of interest (ROIs), allowing for a direct comparison of model prediction accuracy across distinct brain areas.
Figure 9: Multi-model performance distribution across bilateral brain regions, showing correlation coefficients (range: -0.05 to 0.15) for 12 ROIs in 4×3 grid layout

### v3

@@ Introduction
To examine the relationship between hemispheric asymmetry and model performance, Figure 8(c) shows correlations between LH-RH differences and performance metrics. Among the six ROIs analyzed, IFG and MFG lateralization were potentially linked to performance, showing positive relationships (Pearson correlation: IFG: $r=0.54$, $p=0.055$; MFG: $r=0.50$, $p=0.084$). No significant correlations were found for other regions (Pearson correlation: AntTemp: $r=0.02$, $p=0.941$; PostTemp: $r=0.09$, $p=0.770$; AngG: $r=0.09$, $p=0.687$; IFGorb: r=0.06r=-0.06, $p=0.845$). The regulatory role of the prefrontal cortex over distributed brain regions may underlie this phenomenon (Badre and Nee 2018), where its lateralized functional specialization could enhance top-down coordination of cognitive resources, mirroring the efficiency optimization observed in computational models.
This experiment confirms that hemispheric lateralization plays a significant role in the neural representation of language-related processes as captured by LLMs. The observed left-hemisphere dominance in core language areas and right-hemisphere contributions to higher-level cognitive functions align well with existing neurocognitive theories.
Discussion and Conclusion
Understanding how LLMs align with human brain activity during sentence processing offers important insights into the parallels and divergences between artificial and biological language systems (Mahowald et al. 2024; Zhou et al. 2024). Our study contributes to this growing body of work through three core findings: (1) intermediate-layer alignment with neural activity, (2) the effects of instruction tuning and semantic comprehension on brain alignment, and (3) hemispheric asymmetries in neural–LLM correspondence. Below, we discuss each in turn.
Intermediate layers better align with brain activity.
Across all evaluated models, we found that intermediate layers exhibit higher brain–model correlation than final layers, particularly across language-selective brain regions (e.g., IFG, PostTemp). This aligns with previous studies using intracranial EEG and MEG (Mischler et al. 2024; Zhou et al. 2024), and confirms that middle-layer representations in LLMs carry rich linguistic information that most closely mirrors human sentence-level processing.
Our use of fMRI complements these prior modalities by providing broader spatial coverage across regions and hemispheres. These results emphasize the importance of considering layer-wise architecture when interpreting the cognitive fidelity of LLMs, reinforcing the notion that later layers, often optimized for downstream tasks, may abstract away from biologically grounded representations.
Instruction tuning enhances brain alignment and comprehension.
We compared instruction-tuned models with their base versions across both semantic performance and brain activation correlation. Instruction-tuned variants consistently outperformed the base models, with improvements observed in both CSAA metric and neural alignment measures (Figure 6). These gains were statistically significant in performance and showed marginal significance in brain correlation, echoing trends reported in Ren et al. (2025).
Importantly, we shift the interpretive focus away from model size, a dominant narrative in prior work (Bonnasse-Gahot and Pallier 2024), toward semantic comprehension ability as a more meaningful predictor of neural alignment. Our results reveal a scaling law: within the 6.7B–9B parameter range, models with stronger comprehension capabilities (higher CSAA scores) also exhibit significantly higher similarity to human brain activation patterns (Pearson $r=0.601$, $p=0.030$; Figure 7).
This suggests that optimization techniques that improve semantic understanding (such as instruction tuning) not only enhance task performance but also produce representational structures more consistent with human neural activity. Future model evaluation efforts may benefit from incorporating neural alignment as an additional axis of assessment.
Hemispheric asymmetry reflects functional specialization
Our third finding addresses the hemispheric organization of LLM–brain alignment. Consistent with classical neurocognitive theories, we observed left-hemispheric dominance in the inferior frontal gyrus (IFG) and posterior temporal (PostTemp) regions, areas central to syntactic processing and semantic integration (Hu et al. 2023; Vigneau et al. 2006). Conversely, the right hemisphere showed stronger correlations in the middle frontal gyrus (MFG) and anterior temporal (AntTemp), suggesting involvement in metaphor, cross-modal semantics, and cognitive control (Japee et al. 2015; Menon and D’Esposito 2022).
These asymmetries were not only statistically robust (e.g., IFG: $p=0.025$, AntTemp: $p=0.001$; Figure 8b), but also showed positive trends with model performance in specific ROIs. Notably, greater left–right differences in IFG and MFG were associated with higher CSAA scores (Pearson $r=0.54$ and $0.50$, respectively). This suggests that lateralized neural specialization, particularly in prefrontal regions, may support more efficient language representations, paralleling the efficiency gains observed in well-optimized LLMs (Badre and Nee 2018).
Interestingly, other regions (e.g., angular gyrus, IFGorb) showed no significant lateralization–performance correlation, possibly reflecting their roles in bilateral, domain-general processes such as semantic integration or the default mode network (Kuhnke et al. 2023).
Implications and Future Directions.
Together, our findings paint a nuanced picture of how current LLMs reflect, and diverge from, human language systems. Instruction tuning and semantic comprehension capacity, not merely parameter size, are key drivers of neural similarity. Moreover, the correspondence is anatomically and functionally specific: intermediate-layer embeddings align best with neural signals, and this alignment respects known hemispheric specializations in the brain.
These insights raise exciting opportunities for future research. For instance, should training procedures explicitly target alignment with brain data to produce more cognitively plausible models? Could hybrid neuro-symbolic LLMs better reflect distributed and lateralized processing patterns observed in the brain? Our results lay the groundwork for these explorations by bridging computational performance with neurobiological plausibility in future work.